40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3417     Entry time: Thu Aug 12 23:49:04 2010
Author: nancy 
Type: Update 
Category: Environment 
Subject: Laser chiller temp raised 

Since the laser is off, Jenne and I rasied the chiller-chiller (small AC in the Control Room) set point temperature to 73 degree F (from 68F) to save people from shivering.

ELOG V3.1.3-